### v1

# Coupling induced emergent topology in a two-leg fermionic ladder

Abstract
We investigate the ground state properties of spinless fermions on a two leg ladder, by allowing the nearest-neighbour hopping dimerization in one leg and uniform hopping in the other. In the non-interacting limit, we find that, at half-filling, the system exhibits robust topological behavior if the inter-leg hopping is allowed. As a result, a topological phase transition occurs as a function of the inter-leg hopping. When the inter-leg interaction is turned on, the topological phase survives, and we obtain an interaction induced topological phase transition as a function of interaction. Moreover, we show that for both the non-interacting and interacting cases, topological nature of the system survives even by altering the dimerization pattern in the dimerized leg, resulting in two topological phases of opposite character. Finally, we reveal that when uniform interactions are included on all the bonds of the ladder, the topological phase transitions to a symmetry-broken charge-density wave (CDW) phase.
I Introduction
Over the past decades, the notion of topology has emerged as a powerful framework for understanding novel quantum phases of matter that fall beyond the conventional Landau symmetry-breaking paradigm Wilson and Kogut (1974). Unlike conventional phases classified by local order parameters and spontaneous symmetry breaking, topological phases are characterized by global, quantized invariants associated with the wavefunctions of the system. Band topology, in particular, has reshaped the classification of insulating phases in non-interacting systems, giving rise to robust edge or surface states protected by underlying symmetries Hasan and Kane (2010). One such example is the symmetry-protected topological (SPT) phase, which possesses robust boundary states as long as certain symmetries are preserved. The celebrated model that exhibits band topology and an SPT phase is the Su-Schrieffer-Heeger (SSH) model of non-interacting fermions Su et al. (1979), which hosts topologically protected edge states arising from the dimerized hopping in a one-dimensional chain. The SSH model, due to its simplicity, has been extensively studied Asbóth et al. (2016a); Anastasiadis et al. (2022); Li et al. (2014); Pérez-González et al. (2019); Ahmadi et al. (2020), and its topological features have been realized in diverse experimental platforms, including cold atoms, photonic and acoustic lattices, superconducting and electrical circuits, ion traps, and mechanical systems Atala et al. (2013); Nakajima et al. (2016); Lohse et al. (2016); Mukherjee et al. (2017); Lu et al. (2014); Thatcher et al. (2022); Meier et al. (2016); Xie et al. (2019); Kitagawa et al. (2012); Leder et al. (2016); de Léséleuc et al. (2019); Le et al. (2020); Jörg et al. (2024); Liu et al. (2025); Wu et al. (2021); Liu et al. (2022); Lee et al. (2018).
These ideas have now expanded beyond non-interacting limits, opening new avenues for exploring how strong correlations, dimensionality, and symmetry intertwine to stabilize or destabilize topological order in the free fermionic counterpart Rachel (2018). In this context, SPT serves as the simplest playground to study the interplay of these key aspects of quantum condensed matter physics, as these phases are short-range entangled phases Senthil (2015). Due to this, extensive research has been conducted using interacting SSH models to understand how local and non-local interactions affect topological phases and phase transitions in many-body fermionic, bosonic, and spin systems Grusdt et al. (2013); Di Liberto et al. (2016, 2017); Zhou et al. (2023); Li et al. (2013); Taddia et al. (2017); Fraxanet et al. (2022); Julià-Farré et al. (2024a); Mondal et al. (2021); Melo et al. (2023); Márquez et al. (2024); Zhou et al. (2025). An interesting revelation in this context is the interaction induced topology where interactions alone give rise to nontrivial topology in systems that are otherwise topologically trivial in the non-interacting limit Mondal et al. (2022a); Hetényi (2020); Julià-Farré et al. (2022); Zuo et al. (2020); Parida et al. (2025); Julià-Farré et al. (2024b).
On the other hand, interaction effects and dimensionality of the lattices on the topological phases have been investigated by going beyond strictly one-dimensional chains. In particular, ladder geometries or quasi-one dimensional lattices that are intermediate to one and two-dimensional lattices have attracted a great deal of attention, revealing interesting topological properties  Nersesyan (2020); Zhang and Zhou (2017); Li et al. (2017); Padavić et al. (2018); Padhan et al. (2024); Mondal et al. (2023); Feng et al. (2022); Li et al. (2013); Parida et al. (2024); Downing et al. (2024); Ogino et al. (2021, 2022); Liu et al. (2012). The key component in the ladder systems that plays the major role is the inter-layer coupling that decides the fate of the topological properties of individual chains and also imparts novel physical phenomena due to proximity effects Lu and Wang (2007); Shoman et al. (2015); Hsieh et al. (2016); Cheng et al. (2019); Zheng and Hofstetter (2018).
In this work, we study the topological properties of an asymmetric two-leg ladder model composed of spinless fermions, where one leg follows the SSH-type dimerized hopping, and the other is a uniform tight-binding chain. For the non-interacting case, the system becomes topological for any finite dimerization in the dimerized leg which undergoes a transition to a gapped trivial phase as a function of inter-leg (rung) hopping. Interestingly, in the many-body limit, after including the rung interaction, it also drives a topological phase transition in the system as a function of rung interaction. However, when uniform interactions are introduced along both the legs and rungs, this topological phase transition ceases to happen, and the system instead undergoes a transition into a CDW phase.
The paper is organized as follows. In sec. II, we introduce the model under consideration and the computational method used to obtain all the results. In sec. III, we present the non-interacting case, focusing on both regimes of dimerization in the upper leg. Sec. IV discusses the scenario when finite rung interaction is considered. Finally, in sec. VI, we provide a summary of the key results.
II Model
We consider a system of spinless fermions on a two-leg ladder depicted in Fig. 1, where uniform and SSH type dimerized hopping patterns in the lower and the upper legs have been assumed, respectively. The model that describes such a system is given by the Hamiltonian,
$\hat{H}_{M}=\hat{H}+\hat{H}_{V}$ (1)
where,
$\hat{H}$ $=-\bigg(\sum_{i\in\text{odd}}t_{1}\hat{a}_{i}^{\dagger}\hat{a}_{i+1}+\sum_{i\in\text{even}}t_{2}\hat{a}_{i}^{\dagger}\hat{a}_{i+1}+t\sum_{i}\hat{b}_{i}^{\dagger}\hat{b}_{i+1}$
$\;+t_{p}\sum_{i}\hat{a}_{i}^{\dagger}\hat{b}_{i}\bigg)+\text{H.c.}$ (2)
$\hat{H}$ is a single-particle Hamiltonian of non-interacting spinless fermions. $\hat{a}_{i}(\hat{a}_{i}^{\dagger})$ and $\hat{b}_{i}(\hat{b}_{i}^{\dagger})$ are the particle annihilation (creation) operators on leg-a and leg-b, respectively, where $i$ represents the rung index. $t_{1}$ and $t_{2}$ are the alternating hopping amplitudes in the upper leg, $t$ is the uniform hopping strength in the lower leg, and $t_{p}$ represents the strength of the rung (inter-leg) hopping. Throughout all the numerical calculations, we set $t=1$, which sets the energy scale of the system.
Since the single-particle Hamiltonian $\hat{H}$ has a unit cell consisting of four lattice sites, Fourier transforming the Hamiltonian in Eq. (2) yields
$\hat{H}_{k}=\begin{bmatrix}0&\mu_{1}&0&-t_{p}\\ \mu_{1}^{*}&0&-t_{p}&0\\ 0&-t_{p}&0&\mu_{2}\\ -t_{p}&0&\mu_{2}^{*}&0\\ \end{bmatrix}$ (3)
where we have set the lattice constant to unity, $\mu_{1}=-t_{1}-t_{2}e^{-ik}$ and $\mu_{2}=-t(1+e^{ik})$, with $k$ representing the quasi-momentum.
In terms of Pauli matrices, the above Hamiltonian can be written as
$\hat{H}_{k}$ $=$ $-\{[(t_{1}+t)+(t_{2}+t)\text{cos}(k)]\sigma_{x}$ (4)
$+$ $(t_{2}-t)\text{sin}(k)\sigma_{y}\}\otimes\frac{\mathbf{I}}{2}$
- $\{[(t_{1}-t)+(t_{2}-t)\text{cos}(k)]\sigma_{x}$
$+$ $(t_{2}+t)\text{sin}(k)\sigma_{y}\}\otimes\frac{\nu_{z}}{2}-t_{p}\sigma_{x}\otimes\nu_{x}$
where $\nu_{m}$’s and $\sigma_{m}$’s with m=(x,y,z)m=(x,~y,~z), are the Pauli matrices corresponding to the leg of the ladder and sublattice in each leg. $\mathbf{I}$ represents the $2\times 2$ identity matrix.
Figure 1: Schematic of a two-leg ladder with dimerized hopping along the upper leg. The circles denote lattice sites. The alternating hopping amplitudes are $t_{1}$ and $t_{2}$, where the stronger hopping is represented by thick solid lines and the weaker hopping by thin solid lines. The lower leg has uniform hopping, shown by thick solid lines. Panel (a) corresponds to $t_{1}<t_{2}$, while panel (b) corresponds to $t_{1}>t_{2}$.
Upon introducing interactions between the fermions, the above model exhibits characteristics of a strongly correlated Fermi system. For simplicity, we consider only rung-wise repulsive interactions, which can be incorporated as the term $\hat{H}_{V}$ in Eq. (1)
$\hat{H}_{V}$ $=V_{p}\sum_{i}(\hat{n}_{i,a}-\frac{1}{2})(\hat{n}_{i,b}-\frac{1}{2}),$ (5)
where $\hat{n}_{i,a}$ ($\hat{n}_{i,b}$) represents the occupation density in the upper (lower) leg of the ladder.
Figure 2: The left panels in (a-d) represent the single particle dispersion bands for $t_{p}=0.0,~0.8,~\sqrt{2.4},~\text{and}~2.5$ respectively. The right panels in (a-d) represent the corresponding particle filling ($\rho$) plotted against the chemical potential ($\mu$) for the same values of $t_{p}$.
We investigate the many-body ground state properties of the system using the matrix product state (MPS) version of the density matrix renormalization group (DMRG) method Cirac et al. (2021); Schollwöck (2011) under open boundary conditions (OBC). For specific cases, we complement our analysis with exact diagonalization (ED) using periodic boundary conditions (PBC). In the DMRG calculations, the system size is taken up to $L=200$ rungs (=400=400 sites), with a maximum bond dimension of $800$ to minimize truncation errors. A system of size $L=12$ rungs ($24$ sites) is used for ED simulations. Unless otherwise mentioned, all observables are extrapolated to the thermodynamic limit.
III Results
In this section, we first analyze the effect of rung hopping in the non-interacting ladder, and then we investigate the role of rung interaction.
III.1 Non-interacting case
To understand the topological nature of the non-interacting particles, we begin by examining the gap-opening as well as the gap-closing scenarios in the non-interacting picture. Since in fermionic systems, the energy gaps can be fully understood from the underlying band structure, we start by diagonalizing the Hamiltonian given in Eq. (4), and the resulting dispersion relation is
$E^{2}=t_{p}^{2}+\frac{1}{2}(|\mu_{1}|^{2}+|\mu_{2}|^{2})$
$\pm\frac{1}{2}\sqrt{(2t_{p}^{2}+|\mu_{1}|^{2}+|\mu_{2}|^{2})^{2}+4(\mu_{1}\mu_{2}-t_{p}^{2})(t_{p}^{2}-\mu_{1}^{*}\mu_{2}^{*})},$
where, $\mu_{1}$ and $\mu_{2}$ are as defined earlier.
The above dispersion relation suggests that for any finite dimerization in hopping in leg-a, the gap closes at $k=0$ when $t_{p}^{2}=2t(t_{1}+t_{2})$, and at $k=\pi$, when $t_{p}^{2}=0$. At all other points, the gap remains open. We display the band structure for four exemplary values of $t_{p}$: $t_{p}=0$, $0.8$, $\sqrt{2.4}$, and $2.5$ in the left panles of Fig. 2(a–d) respectively by fixing $t_{1}=0.2$ and $t_{2}=1.0$. To compare the band picture with the filling of particles in real space, the filling $\rho=N/2L$ as a function of chemical potential $\mu$ Crépin et al. (2011) calculated under periodic boundary conditions, is plotted in right panels of Fig. 2(a–d), where $N$ is the total number of fermions and $L$ is the number of rungs. The appearance of cusps in the $\rho$$\mu$ curves reflects changes in the number of Fermi points i.e., the number of quasi-momenta where the Fermi energy intersects the bands. The width of each plateau in the $\rho$$\mu$ curve in Fig. 2 indicates the value of the energy gap at the corresponding filling, whereas the shoulder regions around the plateaus indicate the gapless Luttinger liquid. Comparing the two panels in Fig. 2(a), we observe that for $t_{p}=0$, the system remains gapless at all fillings. In contrast, for $t_{p}=0.8$, finite gaps emerge at quarter, half, and three-quarter fillings, as shown in Fig. 2(b). Increasing the value of $t_{p}$ to $\sqrt{2.4}$ closes the gap at half-filling, while the gaps at quarter and three-quarter fillings persist [see Fig. 2(c)]. At $t_{p}=2.5$, gaps reappear again at all three commensurate fillings, such as for quarter, half, and three-quarter fillings, as shown in Fig. 2(d).
With this information in hand about the gap of the system, we investigate topological phase transitions in the subsequent subsection.
III.2 Topological Phase Transition
Since one of the legs of the ladder is considered to be of SSH type, we can have two possible dimerization patterns: (i) topological pattern ($t_{1}<t_{2}$ as in Fig. 1(a)) and (ii) the trivial pattern ($t_{1}>t_{2}$ as in Fig. 1(b)). For the topological (trivial) pattern, we choose $t_{1}=0.2,~t_{2}=1.0$ ($t_{2}=1.0,~t_{2}=0.2$). Note that the hopping in the other leg is uniform throughout and is set to $t=1.0$. This choice of hopping suggests that in the decoupled leg limit, the leg-a can be either trivial or topological insulator whereas the leg-b is always a gapless Luttinger liquid. Our goal is to investigate the influence of the dimerized leg on the system when the two legs are coupled to each other.
Figure 3: (a) and (b) energy spectrum as a function of rung hopping $t_{p}$ when $t_{1}=0.2$, $t_{2}=1.0$ and $t_{1}=1.0$, $t_{2}=0.2$ respectively. (c) and (d) represent the probability density $|{\psi_{i}}|^{2}$, where the circles represent the sites of the ladder and the facecolor of the circle represents the probability density corresponding to the blue circle in (a) and (b) respectively.
We first explore the real space energy spectrum of the system under OBC. In Fig. 3(a) and (b) we depict the energy spectrum as a function of $t_{p}$ for $t_{1}<t_{2}$ and $t_{1}>t_{2}$ configurations, respectively. For both the configurations, the bulk gaps open up as soon as $t_{p}$ becomes finite. Interestingly, for both the topological and trivial dimerization configuration of leg-a, the bulk gap hosts two mid-gap zero-energy states which disappear at a critical $t_{p}=\sqrt{2.4}$ with the closing up of the bulk gap. However, the gap reopens with further increase in $t_{p}$ without the zero energy states. These zero energy state are typical signatures of topological phases. In order to examine that these zero-energy states are edge states, we plot the onsite probability density ($|{\psi_{i}}|^{2}$) corresponding to one of these two states under OBC. As shown in Fig. 3(c) and (d) the states corresponding to the dimerization patterns $t_{1}=0.2,~t_{2}=1.0$ and $t_{1}=1.0,~t_{2}=0.2$ for an exemplary value $t_{p}=1.0$ (dots in Fig. 3(a) and (b)), are localized at the edges of the system. The key distinction between both the cases is that the edge states exist on the upper leg when $t_{1}<t_{2}$ and on the lower leg when $t_{1}>t_{2}$. Interestingly, we observe the lower-leg, which doesn’t have any hopping dimarization, has supported a zero energy edge state. Additionally, the spectra also exhibit isolated energetic mid-gap states at $\rho=1/4$ and $\rho=3/4$ fillings for $t_{1}<t_{2}$ (the red solid lines), which are absent for $t_{1}>t_{2}$. For half-filling, it is evident that there is a pair of localized edge modes existing in the upper leg for the $t_{1}<t_{2}$ case, whereas at filling $\rho=1/4$, the edges of both the legs of the ladder hybridize, resulting in energetic edge states. In contrast, there is also another pair of edge modes at zero energy for $t_{1}>t_{2}$ case as shown in Fig. 3(d), and these are localized in the lower leg of the ladder. From now on, we focus on the zero energy edge states appearing at half-filling to characterize the topological property of our system, as these are the states protected by symmetry.
Figure 4: (a) and (b) illustrate winding number as a function of $t_{p}$ when $t_{1}=0.2$, $t_{2}=1.0$ and $t_{1}=1.0$, $t_{2}=0.2$, respectively.
[…]
$\omega=-\mathfrak{I}\ \ln\prod_{j=0}^{N-1}\text{det}(\langle u_{m,k_{j}}|u_{n,k_{j+1}}\rangle),$ (7)
where, $k_{j}=2\pi j/Na$ is the $j^{\text{th}}$ $k$ point in the Brillouin zone and $|u_{m,k_{j}}\rangle$ denotes the occupied Bloch bands. The determinant is taken as we are dealing with a multiband system. It is well known that a trivial bulk is characterized by $\omega=0$, while a topological bulk possesses a value $\omega=\pi$. We calculate the Berry phase $\omega$ for both $t_{1}<t_{2}$ as well as $t_{1}>t_{2}$ cases by taking the filled bands up to half filling and plot $\omega/\pi$ as a function of $t_{p}$ in Fig. 4(a) and (b), respectively. At $t_{p}=0$, the two legs of the ladder are completely decoupled, with one forming a tight-binding chain for which the gap closes at $k=\pi$, making the Berry phase undefined. For any finite $t_{p}$, however, the system acquires a well-defined Berry phase, which takes the value $\pi$ for both $t_{1}<t_{2}$ and $t_{1}>t_{2}$ for arbitrarily small $t_{p}$, confirming the topological nature of the phases. As $t_{p}$ increases, the bulk gap closes at $t_{p}=\sqrt{2.4}$, beyond which the Berry phase drops to 0, indicating a transition to a trivial phase. This change in $\omega/\pi$ from $1$ to 0 for both $t_{1}<t_{2}$ and $t_{1}>t_{2}$ provides clear evidence of a topological phase transition at $t_{p}=\sqrt{2.4}$. Furthermore, this transition is also accompanied by the disappearance of the zero-energy edge modes present in the topological phase, and this is called the bulk-boundary correspondence in our system.
Based on the above analysis of the appearance of zero-energy edge modes under OBC and the quantized topological invariant $\omega$ under PBC, we now characterize the topological properties of our system as a function of $t_{p}$ for different values of dimerization strength (i.e by fixing $t_{2}=1.0$ and varying $t_{1}$).
Figure 5: The phase diagram in the $t_{p}-t_{1}$ plane is shown for $t_{2}=1.0$. In panel (a), the color bar represents the single-particle excitation gap ($G$) at half-filling, computed for a system of size $L=480$ under OBC. In panel (b), the color bar indicates the winding number ($\omega/\pi$) obtained under PBC. The horizontal lines at $t_{1}=t_{2}=1.0$ indicates the gapless line.
To this end, we compute the energy gap ($G$), defined as the energy difference between the two eigenstates at the middle of the bulk gap (i.e., at half-filling), and is given by $G=E_{L}-E_{L-1}$ under OBC. The plot of $G$ (color coded) as a function of $t_{p}$ and $t_{1}$ while fixing $t_{2}=1.0$ is shown in Fig. 5(a). The blue regions, where $G$ vanishes, indicate the presence of degenerate zero-energy modes at half-filling, whereas the in the gray region, $G$ acquires finite values due to the absence of any degenerate edge states. For a fixed $t_{1}$, $G$ remains zero up to a critical value of $t_{p}$, beyond which it becomes finite, suggesting a topological phase transition. The point at which this topological phase transition occurs agrees well with our analytical boundary at $t_{p}=2t(t_{1}+t_{2})$ (indicated by the red line with circles). Note that at $t_{1}=t_{2}$, the entire bulk is gapless up to $t_{p}\leq 2$ indicated by solid cyan line in Fig. 5(a). In Fig. 5(b) we plot the bulk topological invariant $\omega/\pi$ (the winding number) defined in Eq. (14) for the similar parameters as in Fig. 5(a). The values of $\omega/\pi=1~(0)$ confirms the topological (trivial) regions. This clearly complements the topological to trivial phase transition and the gapless line obtained using the gap $G$ in Fig. 5(a). Note that in the gapless region at $t_{1}=t_{2}=1.0$ and up to $t_{p}\leq 2$, the invariant is ill-defined due to the absence of a bulk energy gap and hence the value is not fixed in these parameter regimes. However, the regions of $t_{1}<t_{2}$ and $t_{1}>t_{2}$ are topologically distinct, which we clarify in the following section.
The above analysis demonstrates that a well-defined topological phase emerges in the system due to $t_{p}$, regardless of the dimerization pattern in leg-a, i.e., whether $t_{1}<t_{2}$ or $t_{1}>t_{2}$. Moreover, an increase in the value of $t_{p}$ favours a topological to trivial phase transition for any fixed dimerization strength. In the following, we will quantify the above topological phase transition occurring at half-filling through a quantity that serves as a dynamic probe for characterizing the above topological phase transitions, and also highlight the phenomenon of reversal of Thouless charge pumping.
III.3 Reversal of Thouless charge pumping
Thouless charge pumping (TCP) Thouless (1983) is a versatile platform to characterize the topological phase in the system through periodic modulation of system parameters. This involves slow and cyclic variation of at least two parameters in the system resulting in a quantized particle transport even in the absence of an external bias. One of the parameters drives the system through a topological phase transition, while an additional staggered potential term breaks the inversion symmetry and maintains a finite bulk gap throughout the cycle. This is crucial for ensuring adiabaticity, which is a key requirement for quantized charge transport. During the adiabatic cycle, the system transitions from a topological phase to a trivial phase and back to the initial topological phase, enclosing the topological phase transition point. This process results in the transport of a quantized number of particles and, the number of particles transported is solely determined by the topological invariant, i.e., the Chern number of the system. TCP has been widely investigated in various theoretical and experimental studies in recent years as a key tool for understanding topological phase transitions Lohse et al. (2016); Nakajima et al. (2016); Citro and Aidelsburger (2023); Hayward et al. (2018); Rice and Mele (1982); Citro and Aidelsburger (2023); Asbóth et al. (2016b); Lin et al. (2020); Ke et al. (2017); Argüello-Luengo et al. (2024); Kuno et al. (2017); Bertok et al. (2022); Mondal et al. (2022b); Hayward et al. (2018); Schweizer et al. (2016); Kraus et al. (2012); Taddia et al. (2017); Wang et al. (2013); Walter et al. (2023); Ke et al. (2020); Padhan et al. (2024); Jürgensen et al. (2023, 2021); Viebahn et al. (2024); Zhu et al. (2024). The Rice-Mele model provides a well-established framework for describing charge pumping protocols.
Figure 6: Instantaneous energy spectrum of the Hamiltonian $\hat{H}_{p}$ as a function of the pumping parameter $\tau$, computed for a system of $L=100$ sites under OBC. Panel (a) corresponds to the parameter set $t_{1}=0.2$, $t_{2}=1.0$, while panel (b) shows results for $t_{1}=1.0$, $t_{2}=0.2$. In both cases, the pumping parameters are fixed at $t_{p_{0}}=1.55$, $\delta_{0}=0.5$, and $\Delta_{0}=0.5$. The highlighted markers in (a) and (b) blue circle, green square, and brown star, indicate specific states whose wavefunction amplitudes are shown in panels (c) and (d), respectively. In panel (c), the top, middle, and bottom subplots show the probability density $|\psi_{i}|^{2}$ plotted against the site index for the states marked in (a). Analogously, panel (d) presents the probability density corresponding to the marked states in (b).
To investigate the topological phase transition occurring as a function of $t_{p}$ at half-filling, we propose a pumping protocol that is governed by the Hamiltonian
$\hat{H}_{p}$ $=-\bigg(\sum_{i\in\text{odd}}t_{1}\hat{a}_{i}^{\dagger}\hat{a}_{i+1}+\sum_{i\in\text{even}}t_{2}\hat{a}_{i}^{\dagger}\hat{a}_{i+1}+t\sum_{i}\hat{b}_{i}^{\dagger}\hat{b}_{i+1}$
$\;+(t_{p_{0}}+\delta(\tau))\sum_{i}\hat{a}_{i}^{\dagger}\hat{b}_{i}\bigg)+\text{H.c.}$
$\;+\sum_{i}\Delta(\tau)((-1)^{i}\hat{a}_{i}^{\dagger}\hat{a}_{i}+(-1)^{i+1}\hat{b}_{i}^{\dagger}\hat{b}_{i}).$ (8)
Here, $\tau$ is the pumping parameter and $\Delta$ ensures that the bulk gap remains open throughout the cycle.
Figure 7: Evolution of polarization $P(\tau)$ as a function of $\tau$ for (a) $t_{1}=0.2,t_{2}=1.0$ and (b) $t_{1}=1.0,t_{2}=0.2$. The plot shows results for three pumping cycles with different values of $t_{p_{0}}$ (=0.8,1.55,2.5=0.8,1.55,2.5), as indicated in the inset. In all cases, $\delta_{0}=0.5~\text{and}~\Delta_{0}=0.5$. The polarization changes are represented by blue squares (cycle 1), green circles (cycle 2), and red diamonds (cycle 3).
($t_{p_{0}},0$) is the origin of the pumping cycle in the $\delta-\Delta$ plane. $\delta(\tau)=\delta_{0}\text{cos}(\tau)$ and $\Delta(\tau)=\Delta_{0}\text{sin}(\tau)$ periodically modulate the rung hopping strength ($t_{p}$) and the staggered potential term, respectively.
To investigate the charge pumping mechanism, we present the instantaneous energy spectrum of the Hamiltonian defined in Eq. (8) for a system of size $L=100$ rungs under OBC. Fig. 6(a) and (b) show the spectra for $t_{1}<t_{2}$ and $t_{1}>t_{2}$, respectively. The cyclic variation of the pumping parameters is represented by cycle-2 (highlighted by a green circle in the insets of Fig. 7(a) and (b)), which encloses the topological phase transition point. In both cases, at half-filling, two in-gap modes cross during the cycle, one ascending from the lower to the upper band and the other descending in the opposite direction after a certain value of the pumping parameter. To demonstrate their edge nature and the evolution during the pump, we plot the probability density of these modes just before and after the half-cycle (i.e. $\tau/2\pi=0.5$), which correspond to the green square and brown star marked points in Figs. 6(a) and (b), respectively. The wave function amplitudes are shown in the middle and lower panels of Figs. 6(c) and (d), respectively. These plots confirm that these modes exchange their position from one edge to another edge across the cycle, realizing topological pumping from one edge site to another edge site of the chain. For $t_{1}<t_{2}$, the edge mode initially localized at the right end of the chain moves to the left after $\tau/2\pi=0.5$, and vice versa for $t_{1}>t_{2}$. This clearly illustrates the reversal of the topological charge pumping direction, governed by the relative dimerization pattern of $t_{1}$ and $t_{2}$. Since the cycle starts from a trivial phase, initially all states near half-filling have finite probability in the bulk, which is shown in the upper panels of Fig. 6(c) and (d) corresponding to the blue circle in Fig. 6(a) and (b), respectively.
Figure 8: (a)The evolution of polarization ($P(\tau)$) as a function of $\tau$ at $t_{p}=2$ cut of the phase diagram shown in Fig. 5 for three different values of $t_{1}$ i.e., $0.5$ (blue squares), $1.0$ (green circles), and $1.5$ (red diamonds). The pumping parameters are set to be $t_{p_{0}}=2.0,\delta_{0}=1.0~~\text{and},~\Delta_{0}=0.6$ (b) The amount of charge pumped ($Q$) by varying $t_{1}$ and fixing the value of $t_{p}$ at $2.0$.
To complement the reversal of the charge pumping phenomenon, we compute the electronic polarization $P$ using the occupied Bloch bands of the Rice-Mele Hamiltonian under PBC as
$P=-\frac{1}{2\pi}\mathfrak{I}\ \ln\prod_{j=0}^{N-1}\text{det}(\langle u_{m,k_{j}}|u_{n,k_{j+1}}\rangle)~(\text{mod}~Qa),$ (9)
where $|u_{m,k_{j}}\rangle$ denotes the occupied Bloch bands of the Hamiltonian
$H_{k}=\begin{bmatrix}\Delta(\tau)&\mu_{1}&0&-t_{p}-\delta(\tau)\\ \mu_{1}^{*}&-\Delta(\tau)&-t_{p}-\delta(\tau)&0\\ 0&-t_{p}-\delta(\tau)&\Delta(\tau)&\mu_{2}\\ -t_{p}-\delta(\tau)&0&\mu_{2}^{*}&-\Delta(\tau)\\ \end{bmatrix}$. (10)
Here, $Q$ denotes the particle charge and $a$ is the lattice constant (set to unity). We calculate the value of $P$ at different values of $\tau$, which we call $P(\tau)$ and the amount of charge pumped in a cycle is given by
$Q=\int_{0}^{2\pi}d\tau~\partial_{\tau}P(\tau).$ (11)
It is evident from Fig. 7(a) that only the cycle 2 results in a smooth variation of $P(\tau)$ from 0 to $1$, corresponding to a quantized charge transfer of $Q=+1$. For the cycle-1, no pumping occurs, while the cycle-3 shows discontinuity in $P(\tau)$, indicating a breakdown of transport. We further explore the scenario by switching the dimerization to $t_{1}=1.0$, $t_{2}=0.2$, while keeping other parameters fixed. In this case, $P(\tau)$ evolves from $1$ to 0 for cycle-2, implying a reversed pumping direction with Q=1Q=-1 [see Fig. 7(b)]. This confirms that the direction of charge transport depends on the dimerization pattern, and altering the dimerization pattern reverses the direction of pumping. To further illustrate the reversal of charge pumping, we fix the rung hopping to $t_{p}=2$ in the phase diagram [Fig. 5], and examine the evolution of polarization $P(\tau)$ for three different values of $t_{1}$: 0.5, 1.0, and 1.5, while keeping $t_{2}=1$ fixed [Fig. 8(a)]. For $t_{1}=0.5$ ($t_{1}<t_{2}$), the polarization smoothly changes from 0 to $1$ over the cycle, indicating a pumped charge of $Q=+1$. At $t_{1}=1.0$, the pumping path intersects the gap-closing point, leading to a breakdown of adiabatic transport and resulting in no net pumping. In contrast, for $t_{1}=1.5$ ($t_{1}>t_{2}$), the polarization evolves from $1$ to 0, corresponding to a pumped charge of Q=1Q=-1, demonstrating a reversal in the direction of charge transport compared to the $t_{1}<t_{2}$ case. To quantify this reversal more generally, we compute the total pumped charge $Q$ as a function of $t_{1}$, keeping all other parameters fixed as in Fig. 8(a). The resulting behavior, shown in Fig. 8(b), clearly shows that the amount of charge pumped changes from $1$ to 1-1 as soon as one crosses the $t_{1}=t_{2}$ point. This behavior captures the transition from positive to negative charge pumping, confirming the reversal of the charge pumping. From the above analysis, it is clear that we find a non-trivial phenomenon of reversal of Thouless charge pumping by periodic modulation of the rung hopping strength. This phenomenon of reversal of TCP also indicates that the phases for $t_{1}<t_{2}$ and $t_{1}>t_{2}$ are topologically distinct. In the following section, we show the many-body properties of the model [Eq. (1)] under consideration.
Figure 9: The phase diagram in the $V_{p}-\mu$ plane keeping $t_{1}=0.2,~t_{2}=1.0,~t_{p}=0.8$ fixed. Dashed lines represent the edge states. All the phase boundaries are extrapolated by taking a maximum system of size $L=100$ rungs.
IV Interaction induced topological phase transition
We now investigate how this topological phase arising in the non-interacting limit evolves upon introducing inter-chain (rung) interactions $V_{p}$. Before analyzing the evolution of topological properties, we first present the bulk phase diagram in the $V_{p}-\mu$ plane, with $t_{1}=0.2,~t_{2}=1.0$ and $t_{p}=0.8$, which is depicted in Fig. 9. The white regions are the gapped phases, and the shaded regions are the gapless phases. The black dashed lines are the edge states, which are extracted by tracking the vanishing up of the single-particle excitation gap defined as $G=\mu_{1}^{+}-\mu_{1}^{-}$, where
$\mu^{+}_{1}=E_{N+1}-E_{N}~~~\text{and}~~~\mu^{-}_{1}=E_{N}-E_{N-1}.$ (12)
The boundaries of the gapped phases (gray solid lines with circles) are identified from the $\rho$ - $\mu$ plots Hikihara et al. (2008); Mishra et al. (2015a, b) for different values of $V_{p}$. Some exemplary behaviour of $\rho$ as a function of $\mu$ are shown in Fig. 10.
Figure 10: The change in density ($\rho$) as a function of change in chemical potential ($\mu$) for different values of $V_{p}$ for a system of $L=100$ rungs. The plateaus at different filling ($\rho=1/4,~1/2,~\text{and}~3/4$) indicate the gap in the system and the length of the plateaus give an estimate of the gaps.
In the $\rho-\mu$ curve, the appearance of plateaus indicates a gap, with the plateau width giving its magnitude. The lower and the upper phase boundaries shown in Fig. 9 are determined from the chemical potential values at the left and right ends of these plateaus in Fig. 10. All the phase boundaries obtained are extrapolated to the thermodynamic limit. It is evident from the phase diagram that there are three gapped phases, at $\rho=1/4,~1/2$ and $3/4$ fillings. The upper (lower) white region is the full (vacuum) state with $\rho=1$ ($\rho=0$).
Figure 11: The figure shows the magnitude of bond energies of all the bonds defined in Eq. (13) and the on-site particle number ($\langle\hat{n}_{i}\rangle$) for different phases corresponding to Fig. 9 with a system consisting of $L=100$ rungs. In the figures, the thickness of a bond is proportional to the respective magnitude of the bond energies, and the face color of the circles represents the values of $\langle\hat{n}_{i}\rangle$. (a) The magnitude of bond energy and the onsite particle density at $1/2$ filling ($N=L$) for $V_{p}=0.2$, (b) $V_{p}=2.8$, (c) represents for $1/4$ filling ($N=L/2$) at $V_{p}=1.4$.
To gain further insight into the nature of these gapped phases, we analyze the many-body ground state by examining the magnitude of the NN bond energies and onsite particle densities at different fillings $\rho=1/4,1/2,\text{and}~3/4$, as shown in the phase diagram in Fig. 9. To do this, we compute the bond energies on the legs ($B_{\alpha,i}$) and rungs ($B_{r,i}$) using the following expressions:
$B_{\alpha,i}=\langle\alpha_{i}^{\dagger}\alpha_{i+1}+\text{H.c.}\rangle,\quad B_{r,i}=\langle a_{i}^{\dagger}b_{i}+\text{H.c.}\rangle$ (13)
where $\alpha=a,b$ denotes the legs, and $r$ denotes the rungs. In Fig. 11, the bond thickness represents the magnitude of the bond energy, while the face color indicates the onsite particle density. The results are obtained for a system of $L=100$ rungs at $V_{p}=0.2$, $1.4$, and $2.8$, and the figure focuses only on the central $10$ rungs to highlight the bulk behavior. At $\rho=1/2$, Fig. 11(a) ($V_{p}=0.2$) shows strong dimerization along the legs, with alternating bond strengths indicating a bond-ordered (BO) phase. In contrast, Fig. 11(b) reveals stronger dimerization along the rungs for $V_{p}=2.8$, confirming a rung-dimerized band insulator (R-BI) phase. The phases at $\rho=1/4$ and $3/4$ are related by particle-hole symmetry, so we present results only for $\rho=1/4$ corresponding to $V_{p}=1.4$. In this case, a single particle is localized within alternate plaquettes and hop only within them, which we call a plaquette-ordered band insulator (PO-BI) phase.
[…]
$\gamma=-\mathfrak{I}\ \ln\prod_{n=0}^{N_{\theta}-1}\langle\psi_{\theta_{n}}|\psi_{\theta_{n+1}}\rangle.$ (14)
where $N_{\theta}$ represents the number of discrete points between 0 to $2\pi$. $\gamma$ becomes 0 in the trivial phase and $\pi$ in the topological phase. We compute the value of $\gamma/\pi$ known as the winding number, and plot it as a function of $V_{p}$ in Fig. 13 (a)(blue squares). The value of $\gamma/\pi$ changes from $1$ to 0 at $V_{p}\sim 2.3$, confirming the topological phase transition as a function of the interaction $V_{p}$.
Figure 12: On-site particle density for (a) $V_{p}=0.2$ and (b) $V_{p}=2.8$. The face color of the circles represents the onsite particle density. To highlight the presence of edge states, only a few rungs near the boundary of a ladder with $L=100$ rungs are displayed.
Figure 13: (a) shows the winding number $\gamma/\pi$ (squares) calculated using ED for a system of $L=12$ rungs and the string order parameter $O_{\text{str}}$ (circles) computed using DMRG for a system of $L=100$ rungs as a function of $V_{p}$ with the other parameters set to $t_{1}=0.2,~t_{2}=1.0,~\text{and}~t_{p}=0.8$. (b) shows the lowest four values of the entanglement spectrum ($\xi_{1},~\xi_{2},~\xi_{3},~\xi_{4}$) as a function of $V$, for a system of size $L=200$ rungs.
The topological phases are often characterized by non-local string correlations den Nijs and Rommelse (1989); Tasaki (1991); Hida (1992); Gibson et al. (2011), which can be captured by the string order parameter defined as
$O_{\text{str}}(r)=-\langle Z_{\text{rung}}(i)e^{i\frac{\pi}{2}\sum_{j=i+1}^{k-1}Z_{\text{rung}}(j)}Z_{\text{rung}}(k)\rangle,$ (15)
where $Z_{\text{rung}}(i)=Z_{a}(i)+Z_{b}(i)$ and $Z_{\alpha}(i)=1-2\alpha_{i}^{\dagger}\alpha_{i}$. Here, $i$ denotes the $i^{th}$ rung of the ladder and r=|ik|r=|i-k| represents the distance between the $i^{th}$ and $k^{th}$ rung. To avoid edge rungs, we have chosen $i=2$ and k=L1k=L-1. We compute $O_{\text{str}}$ for a system of size $L=100$ rungs and plot it along with the winding number in Fig. 13(a) (red circles). A comparison between the behavior of $O_{\text{str}}$ and $\gamma/\pi$, reveals that $O_{\text{str}}$ agrees well with the topological phase transition point identified from the winding number.
To further complement this topological phase transition, we compute the entanglement spectrum of the ground state, which serves as a powerful tool for detecting topological order, as it is related to the number of edge excitations in the system Li and Haldane (2008); Yoshida et al. (2014); Turner et al. (2011); Zhou et al. (2023). The entanglement spectrum is obtained by logarithmically rescaling the eigenvalues of the reduced density matrix corresponding to a specific bipartition in the system. For a bipartition at $l^{th}$ bond, the reduced density matrix is defined as $\rho_{l}=\text{Tr}_{L-l}|\psi\rangle\langle\psi|$, where $|\psi\rangle$ is the ground state wave function. In the MPS representation used in DMRG, the squared singular values $\Lambda_{i}^{2}$ at bond $l$ correspond to the eigenvalues $\lambda_{i}$ of $\rho_{l}$. The entanglement spectrum is then given by
$\xi_{i}~=~-~\text{ln}(\Lambda_{i}^{2}),$ (16)
We compute the entanglement spectrum at the central bond of a system with $L=200$ rungs and plot the four lowest levels $\xi$ as a function of $V_{p}$ in Fig. 13(b). A clear two-fold degeneracy appears in the range $0\leq V_{p}\lesssim 2.3$, which is lifted outside this region. This degeneracy signals the presence of topological order in the ground state for $0\leq V_{p}\lesssim 2.3$.
The above analysis demonstrates that the rung interaction ($V_{p}$) drives a topological phase transition, which in turn enables a Thouless charge pumping mechanism that is modulated by the rung interaction. Additionally, we also observe that the direction of charge pumping reverses for the opposite dimerization configuration in such an interacting system (not shown).
Figure 14: The phase diagram in the $V-\mu$ plane keeping $t_{1}=0.2,~t_{2}=1.0,~t_{p}=0.8$ fixed. Dashed lines represent the edge states. All the phase boundaries are extrapolated by taking a maximum system of size $L=100$ rungs.
V Effect of uniform nearest-neighbor interaction
For completeness, we also examine the case of uniform NN interactions applied equally along both the legs and the rungs, and analyze their effect on the topological phases present in the non-interacting limit.
Figure 15: The figure shows the magnitude of bond energies of all the bonds defined in Eq.(13) and the on-site particle density for different phases corresponding to Fig. 9 with a system consisting of $L=100$ rungs. Only the central $10$ rungs of a system with $L=100$ rungs are shown to focus on the bulk behavior. In the figures, the thickness of a bond is proportional to the respective magnitude of bond energy, and the face color of the circles represents the values of the onsite particle density for (a) $V=0.2$, and (b) $V=2.8$.
The total Hamiltonian in this case is given by
$\hat{H}{{}_{U}}=\hat{H}+\hat{H}_{V_{U}}$ (17)
where $\hat{H}$ is the non-interacting part described earlier, and the interaction term $\hat{H}_{{V_{U}}}$ is defined as
$\hat{H}_{{V_{U}}}$ $=V\sum_{i}\left(\hat{n}_{i,a}-\frac{1}{2}\right)\left(\hat{n}_{i,b}-\frac{1}{2}\right)$
$\quad+V\sum_{i,\alpha}\left(\hat{n}_{i,\alpha}-\frac{1}{2}\right)\left(\hat{n}_{i+1,\alpha}-\frac{1}{2}\right)$ (18)
[…]
VI Conclusion
In this study, we explored the effect of inter-leg coupling in establishing topological phases of spinless fermions on a two-leg ladder. By assuming one of the legs dimerized in nearest-neighbour hopping and the other uniform, we showed that the system as a whole becomes topological as soon as the inter-leg hopping is switched on. However, with increase in the inter-leg hopping, the system undergoes a transition to a trivial phase. We also found that this emergent topological phase in the ladder is independent of the dimerization pattern in the dimerized hopping leg. As a result, we obtained two different topological phases of opposite character. Crucially, we showed that the inter-leg interactions can induce a distinct topological phase transition, revealing an interaction-induced topological phase transition in the system.
We characterized this transition by multiple diagnostics, including the entanglement spectrum, topological invariants, string order parameters and Thouless charge pumping. Our study reveals a non-trivial topological phenomenon involving spinless fermions on a quasi-1D geometry which clearly highlights the intricate interplay between lattice geometry, many-body effects and particle statistics in stabilizing the topological phase transitions in coupled systems. This also opens up possible avenues to explore the fate of these topological phases in bosonic and spinful fermionic many-body systems, in lattices other than ladder geometries, and in the presence of disorder.
VII Acknowledgment
We thank Sanchayan Banerjee for useful discussions. T.M. acknowledges support from Science and Engineering Research Board (SERB), Govt. of India, through project No. MTR/2022/000382 and STR/2022/000023.

### v2

@@
 In the non-interacting limit, we find that, at half-filling, the system exhibits robust topological behavior if the inter-leg hopping is allowed.
+Though depending on the dimerization pattern, the dimerized leg can be either topological or trivial in nature, here we show that by connecting such a leg to a uniform leg through inter-chain coupling, the overall system becomes topological irrespective of the dimerization pattern in the dimerized leg.
 As a result, a topological phase transition occurs as a function of the inter-leg hopping.
-When the inter-leg interaction is turned on, the topological phase survives, and we obtain an interaction induced topological phase transition as a function of interaction.
-Moreover, we show that for both the non-interacting and interacting cases, topological nature of the system survives even by altering the dimerization pattern in the dimerized leg, resulting in two topological phases of opposite character.
+When the inter-leg interaction is turned on, the topological phase survives, and we obtain an interaction induced topological phase transition.
 Finally, we reveal that when uniform interactions are included on all the bonds of the ladder, the topological phase transitions to a symmetry-broken charge-density wave (CDW) phase.
@@
 These ideas have now expanded beyond non-interacting limits, opening new avenues for exploring how strong correlations, dimensionality, and symmetry intertwine to stabilize or destabilize topological order in the free fermionic counterpart Rachel (2018).
-In this context, SPT serves as the simplest playground to study the interplay of these key aspects of quantum condensed matter physics, as these phases are short-range entangled phases Senthil (2015).
+In this context, SPT serves as the simplest playground to study the interplay of these key aspects of quantum condensed matter physics, as these phases are short-range entangled phases Chen et al.
+(2010); Senthil (2015).
 Due to this, extensive research has been conducted using interacting SSH models to understand how local and non-local interactions affect topological phases and phase transitions in many-body fermionic, bosonic, and spin systems Grusdt et al.
@@
 (2024); Zhou et al.
-(2025).
+(2025); Mohamadi and Abouie (2025).
 An interesting revelation in this context is the interaction induced topology where interactions alone give rise to nontrivial topology in systems that are otherwise topologically trivial in the non-interacting limit Mondal et al.
@@
 (2021, 2022); Liu et al.
-(2012).
-The key component in the ladder systems that plays the major role is the inter-layer coupling that decides the fate of the topological properties of individual chains and also imparts novel physical phenomena due to proximity effects Lu and Wang (2007); Shoman et al.
+(2012); Aghtouman and Hosseini (2024); D’Elia and Paiva (2025); Jangjan and Hosseini (2020, 2022).
+The key component in the ladder systems that plays the major role is the inter-layer coupling that decides the fate of the topological properties of individual chains and may impart novel physical phenomena due to proximity effects Lu and Wang (2007); Shoman et al.
 (2015); Hsieh et al.
@@
 In this work, we study the topological properties of an asymmetric two-leg ladder model composed of spinless fermions, where one leg follows the SSH-type dimerized hopping, and the other is a uniform tight-binding chain.
-For the non-interacting case, the system becomes topological for any finite dimerization in the dimerized leg which undergoes a transition to a gapped trivial phase as a function of inter-leg (rung) hopping.
-Interestingly, in the many-body limit, after including the rung interaction, it also drives a topological phase transition in the system as a function of rung interaction.
+For the non-interacting case, the dimerization pattern decides the topological or trivial nature of the SSH-type leg.
+However, once coupled to a uniform tight-binding chain through interchain tunneling, the ladder as a whole becomes topological regardless of the dimerization pattern in the dimerized leg.
+Interestingly, in the many-body limit, including the rung interaction, it also drives a topological phase transition in the system as a function of rung interaction.
 However, when uniform interactions are introduced along both the legs and rungs, this topological phase transition ceases to happen, and the system instead undergoes a transition into a CDW phase.
@@
 In sec.
-III, we present the non-interacting case, focusing on both regimes of dimerization in the upper leg.
+III.1, we present the non-interacting case, focusing on both regimes of dimerization in the SSH-type leg.
 Sec.
-IV discusses the scenario when finite rung interaction is considered.
+III.2 discusses the scenario when interaction between the particles is considered.
 Finally, in sec.
-VI, we provide a summary of the key results.
+IV, we provide a summary of the key results.
 II Model
@@
 The model that describes such a system is given by the Hamiltonian,
-$\hat{H}_{M}=\hat{H}+\hat{H}_{V}$ (1)
+$\hat{H}_{M}=\hat{H}+\hat{H}_{V_{p}}$ (1)
 where,
@@
 $\;+t_{p}\sum_{i}\hat{a}_{i}^{\dagger}\hat{b}_{i}\bigg)+\text{H.c.}$ (2)
-$\hat{H}$ is a single-particle Hamiltonian of non-interacting spinless fermions.
+$\hat{H}$ represents the Hamiltonian of non-interacting spinless fermions.
 $\hat{a}_{i}(\hat{a}_{i}^{\dagger})$ and $\hat{b}_{i}(\hat{b}_{i}^{\dagger})$ are the particle annihilation (creation) operators on leg-a and leg-b, respectively, where $i$ represents the rung index.
@@
 Upon introducing interactions between the fermions, the above model exhibits characteristics of a strongly correlated Fermi system.
-For simplicity, we consider only rung-wise repulsive interactions, which can be incorporated as the term $\hat{H}_{V}$ in Eq.
+For simplicity, we consider only rung-wise repulsive interactions, which can be incorporated as the term $\hat{H}_{V_{p}}$ in Eq.
 (1)
-$\hat{H}_{V}$ $=V_{p}\sum_{i}(\hat{n}_{i,a}-\frac{1}{2})(\hat{n}_{i,b}-\frac{1}{2}),$ (5)
+$\hat{H}_{V_{p}}$ $=V_{p}\sum_{i}(\hat{n}_{i,a}-\frac{1}{2})(\hat{n}_{i,b}-\frac{1}{2}),$ (5)
 where $\hat{n}_{i,a}$ ($\hat{n}_{i,b}$) represents the occupation density in the upper (lower) leg of the ladder.
 Figure 2: The left panels in (a-d) represent the single particle dispersion bands for $t_{p}=0.0,~0.8,~\sqrt{2.4},~\text{and}~2.5$ respectively.
-The right panels in (a-d) represent the corresponding particle filling ($\rho$) plotted against the chemical potential ($\mu$) for the same values of $t_{p}$.
+The right panels in (a-d) represent the corresponding particle filling ($\rho$) plotted against the chemical potential ($\mu$) for the same values of $t_{p}$ with a system of $L=500$ rungs.
+For all the plots, the values of $t_{1}$ and $t_{2}$ are fixed at $0.2$ and $1.0$ respectively.
 We investigate the many-body ground state properties of the system using the matrix product state (MPS) version of the density matrix renormalization group (DMRG) method Cirac et al.
@@
 III Results
-In this section, we first analyze the effect of rung hopping in the non-interacting ladder, and then we investigate the role of rung interaction.
+In this section, we first analyze the noninteracting case, where we investigate the effect of rung hopping, and then we investigate the role of interparticle interaction.
 III.1 Non-interacting case
-To understand the topological nature of the non-interacting particles, we begin by examining the gap-opening as well as the gap-closing scenarios in the non-interacting picture.
+To understand the topological nature of the non-interacting particles, we begin by examining the gap-opening as well as the gap-closing scenarios.
 Since in fermionic systems, the energy gaps can be fully understood from the underlying band structure, we start by diagonalizing the Hamiltonian given in Eq.
@@
 The above dispersion relation suggests that for any finite dimerization in hopping in leg-a, the gap closes at $k=0$ when $t_{p}^{2}=2t(t_{1}+t_{2})$, and at $k=\pi$, when $t_{p}^{2}=0$.
-At all other points, the gap remains open.
+At all other points, the gap remains open at half-filling.
 We display the band structure for four exemplary values of $t_{p}$: $t_{p}=0$, $0.8$, $\sqrt{2.4}$, and $2.5$ in the left panles of Fig.
@@
 To compare the band picture with the filling of particles in real space, the filling $\rho=N/2L$ as a function of chemical potential $\mu$ Crépin et al.
-(2011) calculated under periodic boundary conditions, is plotted in right panels of Fig.
+(2011) calculated under PBC, is plotted in right panels of Fig.
 2(a–d), where $N$ is the total number of fermions and $L$ is the number of rungs.
 The appearance of cusps in the $\rho$$\mu$ curves reflects changes in the number of Fermi points i.e., the number of quasi-momenta where the Fermi energy intersects the bands.
-The width of each plateau in the $\rho$$\mu$ curve in Fig.
-2 indicates the value of the energy gap at the corresponding filling, whereas the shoulder regions around the plateaus indicate the gapless Luttinger liquid.
+The width of each plateau in the $\rho$$\mu$ curve indicates the value of the energy gap at the corresponding filling, whereas the shoulder regions around the plateaus indicate the gapless Luttinger liquid.
 Comparing the two panels in Fig.
 2(a), we observe that for $t_{p}=0$, the system remains gapless at all fillings.
-In contrast, for $t_{p}=0.8$, finite gaps emerge at quarter, half, and three-quarter fillings, as shown in Fig.
+In contrast, for $t_{p}=0.8$, finite gaps emerge at $\rho=1/4,~1/2,~\text{and}~3/4$, as shown in Fig.
 2(b).
-Increasing the value of $t_{p}$ to $\sqrt{2.4}$ closes the gap at half-filling, while the gaps at quarter and three-quarter fillings persist [see Fig.
+Increasing the value of $t_{p}$ to $\sqrt{2.4}$ closes the gap at $\rho=1/2$, while the gaps at $\rho=1/4,~\text{and}~3/4$ persist [see Fig.
 2(c)].
-At $t_{p}=2.5$, gaps reappear again at all three commensurate fillings, such as for quarter, half, and three-quarter fillings, as shown in Fig.
+At $t_{p}=2.5$, gaps reappear again at all three commensurate fillings, such as for $\rho=1/4,~1/2,~\text{and}~3/4$, as shown in Fig.
 2(d).
 With this information in hand about the gap of the system, we investigate topological phase transitions in the subsequent subsection.
-III.2 Topological Phase Transition
+III.1.1 Topological Phase Transition
 Since one of the legs of the ladder is considered to be of SSH type, we can have two possible dimerization patterns: (i) topological pattern ($t_{1}<t_{2}$ as in Fig.
@@
 Note that the hopping in the other leg is uniform throughout and is set to $t=1.0$.
-This choice of hopping suggests that in the decoupled leg limit, the leg-a can be either trivial or topological insulator whereas the leg-b is always a gapless Luttinger liquid.
+This choice of hopping suggests that in the decoupled leg limit, the leg-a can be either trivial or a topological insulator whereas the leg-b is always a gapless Luttinger liquid.
 Our goal is to investigate the influence of the dimerized leg on the system when the two legs are coupled to each other.
 Figure 3: (a) and (b) energy spectrum as a function of rung hopping $t_{p}$ when $t_{1}=0.2$, $t_{2}=1.0$ and $t_{1}=1.0$, $t_{2}=0.2$ respectively.
-(c) and (d) represent the probability density $|{\psi_{i}}|^{2}$, where the circles represent the sites of the ladder and the facecolor of the circle represents the probability density corresponding to the blue circle in (a) and (b) respectively.
+(c) and (d) represent the probability density $|{\psi_{i}}|^{2}$, where the circles represent the sites of the ladder and the facecolor of the circle represents the probability density corresponding to the blue circle in (a) and (b), respectively.
 We first explore the real space energy spectrum of the system under OBC.
@@
 However, the gap reopens with further increase in $t_{p}$ without the zero energy states.
-These zero energy state are typical signatures of topological phases.
-In order to examine that these zero-energy states are edge states, we plot the onsite probability density ($|{\psi_{i}}|^{2}$) corresponding to one of these two states under OBC.
+These zero energy states are typical signatures of topological phases under OBC.
+In order to examine that these zero-energy states are edge states, we plot the onsite probability density ($|{\psi_{i}}|^{2}$) corresponding to one of these two states.
 As shown in Fig.
@@
 Additionally, the spectra also exhibit isolated energetic mid-gap states at $\rho=1/4$ and $\rho=3/4$ fillings for $t_{1}<t_{2}$ (the red solid lines), which are absent for $t_{1}>t_{2}$.
-For half-filling, it is evident that there is a pair of localized edge modes existing in the upper leg for the $t_{1}<t_{2}$ case, whereas at filling $\rho=1/4$, the edges of both the legs of the ladder hybridize, resulting in energetic edge states.
+For $\rho=1/2$, it is evident that there is a pair of localized edge modes existing in the upper leg for the $t_{1}<t_{2}$ case, whereas at $\rho=1/4$, the edges of both the legs of the ladder hybridize (not shown), resulting in energetic edge states.
 In contrast, there is also another pair of edge modes at zero energy for $t_{1}>t_{2}$ case as shown in Fig.
@@
 $\omega=-\mathfrak{I}\ \ln\prod_{j=0}^{N-1}\text{det}(\langle u_{m,k_{j}}|u_{n,k_{j+1}}\rangle),$ (7)
-where, $k_{j}=2\pi j/Na$ is the $j^{\text{th}}$ $k$ point in the Brillouin zone and $|u_{m,k_{j}}\rangle$ denotes the occupied Bloch bands.
+where, $\mathfrak{I}$ denotes the imaginary part, $k_{j}=2\pi j/Na$ is the $j^{\text{th}}$ $k$ point in the Brillouin zone and $|u_{m,k_{j}}\rangle$ denotes the occupied Bloch bands with $m$ and $n$ representing the band indices.
 The determinant is taken as we are dealing with a multiband system.
 It is well known that a trivial bulk is characterized by $\omega=0$, while a topological bulk possesses a value $\omega=\pi$.
-We calculate the Berry phase $\omega$ for both $t_{1}<t_{2}$ as well as $t_{1}>t_{2}$ cases by taking the filled bands up to half filling and plot $\omega/\pi$ as a function of $t_{p}$ in Fig.
+We calculate the Berry phase $\omega$ for both $t_{1}<t_{2}$ as well as $t_{1}>t_{2}$ cases by taking the filled bands up to half filling and plot $\omega/\pi$, called the winding number, as a function of $t_{p}$ in Fig.
 4(a) and (b), respectively.
@@
 Furthermore, this transition is also accompanied by the disappearance of the zero-energy edge modes present in the topological phase, and this is called the bulk-boundary correspondence in our system.
-Based on the above analysis of the appearance of zero-energy edge modes under OBC and the quantized topological invariant $\omega$ under PBC, we now characterize the topological properties of our system as a function of $t_{p}$ for different values of dimerization strength (i.e by fixing $t_{2}=1.0$ and varying $t_{1}$).
+III.1.2 Phase diagram
+Based on the above analysis of the appearance of zero-energy edge modes under OBC and the quantized topological invariant $\omega$ under PBC, we now demonstrate the topological phase transitions in the form of phase diagrams.
+For this purpose, we characterize the topological properties of the system as a function of $t_{p}$ for different values of dimerization strength, that is by fixing $t_{2}=1.0$ and varying $t_{1}$.
 Figure 5: The phase diagram in the $t_{p}-t_{1}$ plane is shown for $t_{2}=1.0$.
 In panel (a), the color bar represents the single-particle excitation gap ($G$) at half-filling, computed for a system of size $L=480$ under OBC.
-In panel (b), the color bar indicates the winding number ($\omega/\pi$) obtained under PBC.
-The horizontal lines at $t_{1}=t_{2}=1.0$ indicates the gapless line.
-To this end, we compute the energy gap ($G$), defined as the energy difference between the two eigenstates at the middle of the bulk gap (i.e., at half-filling), and is given by $G=E_{L}-E_{L-1}$ under OBC.
+In panel (b), the blue regions correspond to a winding number $\omega/\pi=1$ and gray regions to $\omega/\pi=0$ obtained under PBC.
+The horizontal lines at $t_{1}=t_{2}=1.0$ indicate the gapless line.
+To this end, we compute the energy gap ($G$), defined as the energy difference between the two eigenstates at the middle of the bulk gap (i.e., at $\rho=1/2$), and is given by $G=E_{L}-E_{L-1}$ under OBC.
 The plot of $G$ (color coded) as a function of $t_{p}$ and $t_{1}$ while fixing $t_{2}=1.0$ is shown in Fig.
 5(a).
-The blue regions, where $G$ vanishes, indicate the presence of degenerate zero-energy modes at half-filling, whereas the in the gray region, $G$ acquires finite values due to the absence of any degenerate edge states.
+The blue regions, where $G$ vanishes, indicate the presence of degenerate zero-energy modes at $\rho=1/2$, whereas in the gray region, $G$ acquires finite values due to the absence of any degenerate edge states.
 For a fixed $t_{1}$, $G$ remains zero up to a critical value of $t_{p}$, beyond which it becomes finite, suggesting a topological phase transition.
-The point at which this topological phase transition occurs agrees well with our analytical boundary at $t_{p}=2t(t_{1}+t_{2})$ (indicated by the red line with circles).
-Note that at $t_{1}=t_{2}$, the entire bulk is gapless up to $t_{p}\leq 2$ indicated by solid cyan line in Fig.
+The point at which this topological phase transition occurs agrees well with the analytical boundary at $t_{p}=\sqrt{2t(t_{1}+t_{2})}$ (indicated by the red line with circles).
+Note that at $t_{1}=t_{2}$, the entire bulk is gapless up to $t_{p}\leq 2$ indicated by the horizontal solid line in Fig.
 5(a).
@@
 5(b) we plot the bulk topological invariant $\omega/\pi$ (the winding number) defined in Eq.
-(14) for the similar parameters as in Fig.
+(14) for the same parameters as in Fig.
 5(a).
 The values of $\omega/\pi=1~(0)$ confirms the topological (trivial) regions.
-This clearly complements the topological to trivial phase transition and the gapless line obtained using the gap $G$ in Fig.
-5(a).
+This clearly complements the topological to trivial phase transition as a function of $t_{p}$.
 Note that in the gapless region at $t_{1}=t_{2}=1.0$ and up to $t_{p}\leq 2$, the invariant is ill-defined due to the absence of a bulk energy gap and hence the value is not fixed in these parameter regimes.
-However, the regions of $t_{1}<t_{2}$ and $t_{1}>t_{2}$ are topologically distinct, which we clarify in the following section.
 The above analysis demonstrates that a well-defined topological phase emerges in the system due to $t_{p}$, regardless of the dimerization pattern in leg-a, i.e., whether $t_{1}<t_{2}$ or $t_{1}>t_{2}$.
 Moreover, an increase in the value of $t_{p}$ favours a topological to trivial phase transition for any fixed dimerization strength.
-In the following, we will quantify the above topological phase transition occurring at half-filling through a quantity that serves as a dynamic probe for characterizing the above topological phase transitions, and also highlight the phenomenon of reversal of Thouless charge pumping.
-III.3 Reversal of Thouless charge pumping
+In the following, we will quantify the above topological phase transition occurring at $\rho=1/2$ through a quantity that serves as a dynamic probe for characterizing the above topological phase transitions, and also highlight the phenomenon of reversal of Thouless charge pumping.
+III.1.3 Reversal of Thouless charge pumping
 Thouless charge pumping (TCP) Thouless (1983) is a versatile platform to characterize the topological phase in the system through periodic modulation of system parameters.
@@
 This is crucial for ensuring adiabaticity, which is a key requirement for quantized charge transport.
-During the adiabatic cycle, the system transitions from a topological phase to a trivial phase and back to the initial topological phase, enclosing the topological phase transition point.
+During the adiabatic cycle, the system traverses from a topological phase to a trivial phase and back to the initial topological phase, enclosing the topological phase transition point.
 This process results in the transport of a quantized number of particles and, the number of particles transported is solely determined by the topological invariant, i.e., the Chern number of the system.
@@
 The Rice-Mele model provides a well-established framework for describing charge pumping protocols.
-Figure 6: Instantaneous energy spectrum of the Hamiltonian $\hat{H}_{p}$ as a function of the pumping parameter $\tau$, computed for a system of $L=100$ sites under OBC.
+Figure 6: Instantaneous energy spectrum of the Hamiltonian $\hat{H}_{p}$ as a function of the pumping parameter $\tau$, computed for a system of $L=50$ rungs under OBC.
 Panel (a) corresponds to the parameter set $t_{1}=0.2$, $t_{2}=1.0$, while panel (b) shows results for $t_{1}=1.0$, $t_{2}=0.2$.
 In both cases, the pumping parameters are fixed at $t_{p_{0}}=1.55$, $\delta_{0}=0.5$, and $\Delta_{0}=0.5$.
-The highlighted markers in (a) and (b) blue circle, green square, and brown star, indicate specific states whose wavefunction amplitudes are shown in panels (c) and (d), respectively.
+The highlighted markers in (a) and (b) blue circle, green square, and brown star, indicate specific states whose probability amplitudes are shown in panels (c) and (d), respectively.
 In panel (c), the top, middle, and bottom subplots show the probability density $|\psi_{i}|^{2}$ plotted against the site index for the states marked in (a).
 Analogously, panel (d) presents the probability density corresponding to the marked states in (b).
-To investigate the topological phase transition occurring as a function of $t_{p}$ at half-filling, we propose a pumping protocol that is governed by the Hamiltonian
+To investigate the topological phase transition occurring as a function of $t_{p}$ at $\rho=1/2$, we propose a pumping protocol that is governed by the Hamiltonian
 $\hat{H}_{p}$ $=-\bigg(\sum_{i\in\text{odd}}t_{1}\hat{a}_{i}^{\dagger}\hat{a}_{i+1}+\sum_{i\in\text{even}}t_{2}\hat{a}_{i}^{\dagger}\hat{a}_{i+1}+t\sum_{i}\hat{b}_{i}^{\dagger}\hat{b}_{i+1}$
 $\;+(t_{p_{0}}+\delta(\tau))\sum_{i}\hat{a}_{i}^{\dagger}\hat{b}_{i}\bigg)+\text{H.c.}$
-$\;+\sum_{i}\Delta(\tau)((-1)^{i}\hat{a}_{i}^{\dagger}\hat{a}_{i}+(-1)^{i+1}\hat{b}_{i}^{\dagger}\hat{b}_{i}).$ (8)
+$\;+\sum_{i}\Delta(\tau)((-1)^{i+1}\hat{a}_{i}^{\dagger}\hat{a}_{i}+(-1)^{i}\hat{b}_{i}^{\dagger}\hat{b}_{i}).$ (8)
 Here, $\tau$ is the pumping parameter and $\Delta$ ensures that the bulk gap remains open throughout the cycle.
@@
 7(a) and (b)), which encloses the topological phase transition point.
-In both cases, at half-filling, two in-gap modes cross during the cycle, one ascending from the lower to the upper band and the other descending in the opposite direction after a certain value of the pumping parameter.
+In both cases, at $\rho=1/2$, two in-gap modes cross during the cycle, one ascending from the lower to the upper band and the other descending in the opposite direction after a certain value of the pumping parameter.
 To demonstrate their edge nature and the evolution during the pump, we plot the probability density of these modes just before and after the half-cycle (i.e.
@@
 6(a) and (b), respectively.
-The wave function amplitudes are shown in the middle and lower panels of Figs.
+The probability amplitudes are shown in the middle and lower panels of Figs.
 6(c) and (d), respectively.
@@
 6(a) and (b), respectively.
-Figure 8: (a)The evolution of polarization ($P(\tau)$) as a function of $\tau$ at $t_{p}=2$ cut of the phase diagram shown in Fig.
-5 for three different values of $t_{1}$ i.e., $0.5$ (blue squares), $1.0$ (green circles), and $1.5$ (red diamonds).
-The pumping parameters are set to be $t_{p_{0}}=2.0,\delta_{0}=1.0~~\text{and},~\Delta_{0}=0.6$ (b) The amount of charge pumped ($Q$) by varying $t_{1}$ and fixing the value of $t_{p}$ at $2.0$.
+Figure 8: (a)The evolution of polarization ($P(\tau)$) as a function of $\tau$ at $t_{p_{0}}=2$ for three different values of $t_{1}$ i.e., $0.5$ (blue squares), $1.0$ (green circles), and $1.5$ (red diamonds), keeping $t_{2}=1$ (Refer Phase diagram shown in Fig.
+5).
+The pumping parameters are set to be $t_{p_{0}}=2.0,\delta_{0}=1.0~~\text{and},~\Delta_{0}=0.6$ (b) The amount of charge pumped ($Q$) by varying $t_{1}$ with fixed value of $t_{p_{0}}=2.0$.
 To complement the reversal of the charge pumping phenomenon, we compute the electronic polarization $P$ using the occupied Bloch bands of the Rice-Mele Hamiltonian under PBC as
@@
 This behavior captures the transition from positive to negative charge pumping, confirming the reversal of the charge pumping.
-From the above analysis, it is clear that we find a non-trivial phenomenon of reversal of Thouless charge pumping by periodic modulation of the rung hopping strength.
+From the above analysis, it is clear that we find the phenomenon of reversal of Thouless charge pumping by periodic modulation of the rung hopping strength.
 This phenomenon of reversal of TCP also indicates that the phases for $t_{1}<t_{2}$ and $t_{1}>t_{2}$ are topologically distinct.
-In the following section, we show the many-body properties of the model [Eq.
-(1)] under consideration.
+Note that such topological phases do not exist for hardcore bosonic ladder, and it is solely due to fermionic statistics.
+III.2 Interacting case
+We now investigate how this topological phase arising in the non-interacting limit evolves upon introducing interactions in the system.
+We first explore the effect of inter-chain (rung) repulsive interactions $V_{p}$ and then the effect of repulsive interactions along all the bonds of the system.
 Figure 9: The phase diagram in the $V_{p}-\mu$ plane keeping $t_{1}=0.2,~t_{2}=1.0,~t_{p}=0.8$ fixed.
@@
 All the phase boundaries are extrapolated by taking a maximum system of size $L=100$ rungs.
-IV Interaction induced topological phase transition
-We now investigate how this topological phase arising in the non-interacting limit evolves upon introducing inter-chain (rung) interactions $V_{p}$.
+III.2.1 Effect of inter-chain interaction
 Before analyzing the evolution of topological properties, we first present the bulk phase diagram in the $V_{p}-\mu$ plane, with $t_{1}=0.2,~t_{2}=1.0$ and $t_{p}=0.8$, which is depicted in Fig.
@@
 10.
-Figure 10: The change in density ($\rho$) as a function of change in chemical potential ($\mu$) for different values of $V_{p}$ for a system of $L=100$ rungs.
-The plateaus at different filling ($\rho=1/4,~1/2,~\text{and}~3/4$) indicate the gap in the system and the length of the plateaus give an estimate of the gaps.
 In the $\rho-\mu$ curve, the appearance of plateaus indicates a gap, with the plateau width giving its magnitude.
@@
 The upper (lower) white region is the full (vacuum) state with $\rho=1$ ($\rho=0$).
-Figure 11: The figure shows the magnitude of bond energies of all the bonds defined in Eq.
-(13) and the on-site particle number ($\langle\hat{n}_{i}\rangle$) for different phases corresponding to Fig.
-9 with a system consisting of $L=100$ rungs.
-In the figures, the thickness of a bond is proportional to the respective magnitude of the bond energies, and the face color of the circles represents the values of $\langle\hat{n}_{i}\rangle$.
-(a) The magnitude of bond energy and the onsite particle density at $1/2$ filling ($N=L$) for $V_{p}=0.2$, (b) $V_{p}=2.8$, (c) represents for $1/4$ filling ($N=L/2$) at $V_{p}=1.4$.
-To gain further insight into the nature of these gapped phases, we analyze the many-body ground state by examining the magnitude of the NN bond energies and onsite particle densities at different fillings $\rho=1/4,1/2,\text{and}~3/4$, as shown in the phase diagram in Fig.
+To gain further insight into the nature of these gapped phases, we analyze the many-body ground state by examining the magnitude of the NN bond energies and onsite particle densities at different fillings $\rho=1/4,1/2,~\text{and}~3/4$, as shown in the phase diagram in Fig.
 9.
-To do this, we compute the bond energies on the legs ($B_{\alpha,i}$) and rungs ($B_{r,i}$) using the following expressions:
-$B_{\alpha,i}=\langle\alpha_{i}^{\dagger}\alpha_{i+1}+\text{H.c.}\rangle,\quad B_{r,i}=\langle a_{i}^{\dagger}b_{i}+\text{H.c.}\rangle$ (13)
+To do this, we compute the bond energies on the legs ($B_{i,\alpha}$) and rungs ($B_{i,r}$) using the following expressions:
+$B_{i,\alpha}=\langle\hat{\alpha}_{i}^{\dagger}\hat{\alpha}_{i+1}+\text{H.c.}\rangle,\quad B_{i,r}=\langle\hat{a}_{i}^{\dagger}\hat{b}_{i}+\text{H.c.}\rangle$ (13)
 where $\alpha=a,b$ denotes the legs, and $r$ denotes the rungs.
@@
 In this case, a single particle is localized within alternate plaquettes and hop only within them, which we call a plaquette-ordered band insulator (PO-BI) phase.
-[…]
+Figure 10: The change in particle filling ($\rho$) as a function of change in chemical potential ($\mu$) for different values of $V_{p}$ for a system of $L=100$ rungs keeping $t_{1}=0.2,~t_{2}=1.0,~\text{and}~t_{p}=0.8$ fixed.
+The plateaus at different filling ($\rho=1/4,~1/2,~\text{and}~3/4$) indicate the gap in the system and the length of the plateaus give an estimate of the gaps.
+Building on this understanding of the bulk phases, we now examine how the topological characteristics identified in the non-interacting limit evolve as a function of rung interaction.
+Note that the black dashed lines in Fig.
+9 are the edge states, which also existed in the non-interacting limit (see Fig.
+3).
+We obtain that these edge states also survive in the presence of interaction and at $\rho=1/2$, with an increase in $V_{p}$, a gap-closing phase transition occurs to a trivial phase at $V_{p}\sim 2.3$, which we call the interaction-induced topological phase transition.
+To show that these states exist at the edges in the presence of finite interaction as well, we plot the on-site particle density of both the legs as a function of the rung index for $V_{p}=0.2$ in Fig.
+12(a).
+For comparison, we also show the onsite particle density $\langle\hat{n}_{i,\alpha}\ranglecorresponding$ to the trivial phase, i.e., at $V_{p}=2.8$, which does not have any edge states as shown in Fig.
+12(b).
+The clear asymmetry in density observed in both the edge sites for $V_{p}=0.2$ confirms the edge states.
+In contrast, for $V_{p}=2.8$, the density remains uniform throughout, confirming that there are no edge states in the system.
 […]
@@
 The value of $\gamma/\pi$ changes from $1$ to 0 at $V_{p}\sim 2.3$, confirming the topological phase transition as a function of the interaction $V_{p}$.
-Figure 12: On-site particle density for (a) $V_{p}=0.2$ and (b) $V_{p}=2.8$.
-The face color of the circles represents the onsite particle density.
+Figure 11: The figure shows the magnitude of bond energies of all the bonds defined in Eq.
+(13) and the on-site particle density for different phases corresponding to Fig.
+9 with a system consisting of $L=100$ rungs keeping $t_{1}=0.2,~t_{2}=1.0,~\text{and}~t_{p}=0.8$ fixed.
+In the figures, the thickness of a bond is proportional to the respective magnitude of the bond energies, and the face color of the circles represents the values of $\langle\hat{n}_{i,\alpha}\rangle$.
+(a) The magnitude of bond energy and the onsite particle density at $1/2$ filling ($N=L$) for $V_{p}=0.2$, (b) $V_{p}=2.8$, and (c) represents for $1/4$ filling ($N=L/2$) at $V_{p}=1.4$.
+Figure 12: On-site particle density for (a) $V_{p}=0.2$ and (b) $V_{p}=2.8$ keeping $t_{1}=0.2,~t_{2}=1.0,~\text{and}~t_{p}=0.8$ fixed.
+The face color of the circles represents the onsite particle density $\langle\hat{n}_{i,\alpha}\rangle$.
 To highlight the presence of edge states, only a few rungs near the boundary of a ladder with $L=100$ rungs are displayed.
 Figure 13: (a) shows the winding number $\gamma/\pi$ (squares) calculated using ED for a system of $L=12$ rungs and the string order parameter $O_{\text{str}}$ (circles) computed using DMRG for a system of $L=100$ rungs as a function of $V_{p}$ with the other parameters set to $t_{1}=0.2,~t_{2}=1.0,~\text{and}~t_{p}=0.8$.
-(b) shows the lowest four values of the entanglement spectrum ($\xi_{1},~\xi_{2},~\xi_{3},~\xi_{4}$) as a function of $V$, for a system of size $L=200$ rungs.
+(b) shows the lowest four values of the entanglement spectrum ($\xi_{1},~\xi_{2},~\xi_{3},~\xi_{4}$) as a function of $V_{p}$, for a system of size $L=200$ rungs.
 The topological phases are often characterized by non-local string correlations den Nijs and Rommelse (1989); Tasaki (1991); Hida (1992); Gibson et al.
 (2011), which can be captured by the string order parameter defined as
-$O_{\text{str}}(r)=-\langle Z_{\text{rung}}(i)e^{i\frac{\pi}{2}\sum_{j=i+1}^{k-1}Z_{\text{rung}}(j)}Z_{\text{rung}}(k)\rangle,$ (15)
-where $Z_{\text{rung}}(i)=Z_{a}(i)+Z_{b}(i)$ and $Z_{\alpha}(i)=1-2\alpha_{i}^{\dagger}\alpha_{i}$.
+$O_{\text{str}}(r)=-\langle\hat{Z}_{\text{rung}}(i)e^{i\frac{\pi}{2}\sum_{j=i+1}^{k-1}\hat{Z}_{\text{rung}}(j)}\hat{Z}_{\text{rung}}(k)\rangle,$ (15)
+where $\hat{Z}_{\text{rung}}(i)=\hat{Z}_{a}(i)+\hat{Z}_{b}(i)$ and $\hat{Z}_{\alpha}(i)=1-2\hat{\alpha}_{i}^{\dagger}\hat{\alpha}_{i}$.
 Here, $i$ denotes the $i^{th}$ rung of the ladder and r=|ik|r=|i-k| represents the distance between the $i^{th}$ and $k^{th}$ rung.
@@
 All the phase boundaries are extrapolated by taking a maximum system of size $L=100$ rungs.
-V Effect of uniform nearest-neighbor interaction
+III.2.2 Effect of uniform interaction
 For completeness, we also examine the case of uniform NN interactions applied equally along both the legs and the rungs, and analyze their effect on the topological phases present in the non-interacting limit.
-Figure 15: The figure shows the magnitude of bond energies of all the bonds defined in Eq.(13) and the on-site particle density for different phases corresponding to Fig.
-9 with a system consisting of $L=100$ rungs.
-Only the central $10$ rungs of a system with $L=100$ rungs are shown to focus on the bulk behavior.
-In the figures, the thickness of a bond is proportional to the respective magnitude of bond energy, and the face color of the circles represents the values of the onsite particle density for (a) $V=0.2$, and (b) $V=2.8$.
 The total Hamiltonian in this case is given by
-$\hat{H}{{}_{U}}=\hat{H}+\hat{H}_{V_{U}}$ (17)
-where $\hat{H}$ is the non-interacting part described earlier, and the interaction term $\hat{H}_{{V_{U}}}$ is defined as
-$\hat{H}_{{V_{U}}}$ $=V\sum_{i}\left(\hat{n}_{i,a}-\frac{1}{2}\right)\left(\hat{n}_{i,b}-\frac{1}{2}\right)$
+$\hat{H}{{}_{U}}=\hat{H}+\hat{H}_{V}$ (17)
+where $\hat{H}$ is the non-interacting part described earlier (Eq. (2)), and the interaction term $\hat{H}_{{V}}$ is defined as
+$\hat{H}_{V}$ $=V\sum_{i}\left(\hat{n}_{i,a}-\frac{1}{2}\right)\left(\hat{n}_{i,b}-\frac{1}{2}\right)$
 $\quad+V\sum_{i,\alpha}\left(\hat{n}_{i,\alpha}-\frac{1}{2}\right)\left(\hat{n}_{i+1,\alpha}-\frac{1}{2}\right)$ (18)
 […]
-VI Conclusion
-In this study, we explored the effect of inter-leg coupling in establishing topological phases of spinless fermions on a two-leg ladder.
-By assuming one of the legs dimerized in nearest-neighbour hopping and the other uniform, we showed that the system as a whole becomes topological as soon as the inter-leg hopping is switched on.
-However, with increase in the inter-leg hopping, the system undergoes a transition to a trivial phase.
-We also found that this emergent topological phase in the ladder is independent of the dimerization pattern in the dimerized hopping leg.
-As a result, we obtained two different topological phases of opposite character.
-Crucially, we showed that the inter-leg interactions can induce a distinct topological phase transition, revealing an interaction-induced topological phase transition in the system.
-We characterized this transition by multiple diagnostics, including the entanglement spectrum, topological invariants, string order parameters and Thouless charge pumping.
-Our study reveals a non-trivial topological phenomenon involving spinless fermions on a quasi-1D geometry which clearly highlights the intricate interplay between lattice geometry, many-body effects and particle statistics in stabilizing the topological phase transitions in coupled systems.
+Figure 15: The figure shows the magnitude of bond energies of all the bonds defined in Eq.
+(13) and the on-site particle density for different phases corresponding to Fig.
+14 with a system consisting of $L=100$ rungs keeping $t_{1}=0.2,~t_{2}=1.0,~t_{p}=0.8$ fixed.
+Only the central $10$ rungs of a system with $L=100$ rungs are shown to focus on the bulk behavior.
+In the figures, the thickness of a bond is proportional to the respective magnitude of bond energy, and the face color of the circles represents the values of $\langle\hat{n}_{i,\alpha}\rangle$ for (a) $V=0.2$, and (b) $V=2.8$.
+IV Conclusion
+In this study, we have explored the effect of inter-leg coupling in establishing topological phases and phase transitions of spinless fermions on a two-leg ladder.
+By assuming that one of the legs possessing nearest-neighbour hopping dimerization and the other leg a uniform chain, we have shown that the entire ladder exhibits topological character irrespective of the dimerization pattern in the dimerized leg.
+As a result of this, we obtained two different topological phases of opposite character by changing the dimerization pattern of the leg possessing dimerized hopping.
+However, stronger inter-leg hopping makes the entire system a trivial insulator.
+Additionally, we showed that the inter-leg interactions can induce a distinct topological phase transition, revealing an interaction-induced topological phase transition in the system.
+We characterized these transitions by multiple diagnostics, including the entanglement spectrum, topological invariants, string order parameters and Thouless charge pumping.
+Our study reveals topological phenomena involving spinless fermions on a quasi-1D geometry, which clearly highlights the intricate interplay between lattice geometry, many-body effects, and particle statistics in stabilizing the topological phase transitions in coupled systems.
 This also opens up possible avenues to explore the fate of these topological phases in bosonic and spinful fermionic many-body systems, in lattices other than ladder geometries, and in the presence of disorder.
-VII Acknowledgment
+V Acknowledgment
 We thank Sanchayan Banerjee for useful discussions.
@@
 MTR/2022/000382 and STR/2022/000023.
+VI Data availability
+The data that support the findings of this article are openly available Parida et al. .